# Creativity and Machine Learning: A Survey

### v4

@@ 1. Introduction
The connection between creativity and machines is as old as computer science itself. In 1842, Lady Lovelace, an English mathematician and writer recognized by many as the first computer programmer, issued the so-called “Lovelace’s objection” (Turing, 1950): the Analytical Engine (the digital programmable machine proposed by Charles Babbage (Babbage, 1864)) “has no pretensions to originate anything. It can do whatever we know how to order it to perform” (Menabrea and Lovelace, 1843). Indeed, in the following centuries, several projects and studies concerning the design of machines able to “originate something” have been carried out (Cohen, 1988; Cope, 1989; Hofstadter and Mitchell, 1994; Langley et al., 1987; Meehan, 1977; Racter, 1984). This has led to the emergence of a specialized field in computer science, namely Computational Creativity (Cardoso et al., 2009), which concerns the study of the relationship between creativity and artificial systems (Colton and Wiggins, 2012; Wiggins, 2006).
@@
Related surveys. We now provide an overview of other surveys in areas related to the present work. For readers interested in a survey in deep generative models, we recommend (Bond-Taylor et al., 2021; Harshvardhan et al., 2020); for an analysis of the state of the art in evaluation in computational creativity (Lamb et al., 2018) is an essential reading; for a review on AI and creativity in general, we recommend (Rowe and Partridge, 1993); for a practical view of generative deep learning, we suggest (Foster, 2019); finally, for an in-depth examination of artistic AI works (also in human-computer co-creations (Guzdial and Riedl, 2019)), (Miller, 2019) is a very comprehensive source of information.
@@ 2. Defining Creativity
For this reason, we consider Boden’s three criteria for studying machine creativity, defined as “the ability to come up with ideas or artifacts that are new, surprising and valuable (Boden, 2003). In particular, value refers to utility, performance, and attractiveness (Maher, 2010); it is related with both the quality of the production, and its acceptance by the society. Novelty refers to the dissimilarity between the produced artifact and other examples in its class (Ritchie, 2007). Finally, surprise refers to the degree to which a stimulus disagrees with expectation (Berlyne, 1971). We underpin our analysis on Boden’s three criteria since they have been widely adopted. Boden also suggests that three forms of creativity can be identified (Boden, 2003) to describe how a novel and surprising product is obtained. The three forms of creativity are combinatorial, exploratory and transformational, ordered by increasing rarity and produced surprise. Combinatorial creativity is about making unfamiliar combinations of familiar ideas, e.g., analogies in textual forms or collage in the visual arts. Exploratory creativity involves the exploration of the conceptual space defined by the cultural context considered, e.g., inventing a new type of cut for fries. Transformational creativity involves changing that space in a way that allows new and previously inconceivable thoughts to become possible, as it has been for free verse in poetry or abstract painting in art.
Finally, it is worth noting that Boden also identified four different questions that emerge when studying computational creativity. These are referred to as Lovelace questions, because many people would respond to them by using Lovelace’s objection. The first question is whether computational ideas can help us understand how human creativity works. The second is whether computers could ever do things which at least appear to be creative. The third is whether a computer could ever appear to recognize creativity. Finally, the fourth is whether computers themselves could ever really be creative, i.e., with the originality of the products only deriving from the machine itself (Boden, 2003). While the first one is studied in Boden’s work, we will provide the reader with an overview of the techniques that can possibly be used to answer the second (Section 3) and the third (Section 4). With respect to the fourth, Boden states that “it is not a scientific question as the others are, but in part a philosophical worry about “meaning” and in part a disguised request for a moral political decision”. We agree with this position, but we hope that our survey will provide the reader with elements for answering the fourth one as well.
@@ 3. Generative Models
At a first glance, this definition appears to be incompatible with those presented in Section 2. Indeed, mimicry is the opposite of novelty. However, what a generative model should aim at mimicking is the underlying distribution representative of the artifacts, and not the specific artifacts themselves; in other words, it should aim at learning the conceptual space defined by the cultural context considered. A generative model can be said to exhibit combinatorial creativity if it can sample new and valuable works that are combinations of real data and to exploratory creativity if the works actually differ from real ones. Vice versa, transformational creativity emerges if and only if the distribution for sampling diverges in some way from the underlying one (e.g., due to a different training process, by altering the distribution after learning, or by changing the sampling technique). In summary, what matters is how the space of solutions is learned and how artifacts are sampled from it.
In this section, we aim at studying the level of creativity of existing generative deep learning models. Following the discussion above, we analyze how the models learn their spaces of solutions and how the observations are generated from them. A new generative deep learning taxonomy is then introduced based on the different training and sampling techniques at the basis of each method. Figure 1 provides a summary of the seven generative classes considered in this survey.
@@
Since our focus is on machine creativity, we do not discuss the implementation details of each class of methods. We instead present the core concepts at the basis of each class of methods; some relevant examples of models; potential applications; and a critical discussion evaluating the level of machine creativity considering the definitions above.
@@ 3.1.1. Core Concepts
A Variational Auto-Encoder (VAE) (Kingma and Welling, 2014; Rezende et al., 2014) is a learning architecture composed by two models: an encoder (or recognition model) and a decoder (or generative model). The former compresses high-dimensional input data into a latent space, i.e., a lower-dimensional space whose features are not directly observable, yet provide a meaningful representation. The latter decompresses the representation vector back to the original domain (Foster, 2019). Classic autoencoders directly learn to represent each input in a latent representation vector. Conversely, VAEs learn a (Gaussian) distribution over the possible values of the latent representation, i.e., the encoder learns the mean and the (log of the) variance of the distribution.
@@
The mathematical derivation of the whole loss has its roots in variational inference (Jordan et al., 1999). Indeed, VAEs can be seen as an efficient and stochastic variational inference method, in which neural networks (NNs) and stochastic gradient descent are used to learn an approximation (i.e., the encoder) of the true posterior (Ganguly and Earp, 2021).
In VAEs, similar high-dimensional data are mapped to close distributions. This makes possible to sample a random point $\mathbf{z}$ from the latent space, and still obtain a comprehensible reconstruction (Foster, 2019). On the other hand, VAE tends to produce blurred images (Zhao et al., 2017). It may also happen that high-density regions under the prior have a low density under the approximate posterior, i.e., these regions are not decoded to data-like samples (Aneja et al., 2021). Finally, the objective can lead to overly simplified representations without using the entire capacity, obtaining only a sub-optimal generative model (Burda et al., 2016).
@@ 3.1.2. Examples of Models
Several models based on VAEs have been proposed (Kingma and Welling, 2019) in the recent years. In the following, we focus on those relevant for our discussion on machine creativity.
In $\beta$-VAE (Higgins et al., 2017), a parameter $\beta$ is used to scale the magnitude of the regularization loss, which allows a better disentanglement of the latent space (Burgess et al., 2018). Another example is VAE-GAN (Larsen et al., 2016), which merges VAE and Generative Adversarial Networks (GAN; see Section 3.2) (Goodfellow et al., 2014). This is done by treating the decoder as the generator of the GAN, thus training it by means of the GAN loss function. This leads to the generation of substantially less blurred images. Similarly, Adversarially Learned Inference (ALI) (Dumoulin et al., 2017) merges VAE and GAN by asking the discriminator to distinguish between pairs of real data (and their latent representations) and pairs of sampled representations and synthetic data. Instead, Adversarial Autoencoders (AAE) (Makhzani et al., 2016) substitute the regularization loss with a discriminative signal, where the discriminator has to distinguish between random latent samples and encoded latent vectors. Another way to address the problem of “sample blurriness” is with PixelVAE (Gulrajani et al., 2017), where the autoregressive PixelCNN (Van Den Oord et al., 2016b, c) is used as the decoder. In (Bowman et al., 2016), to deal with sequential data such as texts, where generation requires more steps, the encoder learns to produce a latent representation of a sentence, while the recurrent neural network RNN-based decoder learns to reproduce it word after word. However, VAE can also generate text by means of convolution and deconvolution (Semeniuta et al., 2017).
To solve the problem of low-density regions, the authors of (Aneja et al., 2021) propose an energy-based model called noise contrastive prior (NCP), trained by contrasting samples from the aggregate posterior to samples from a base prior. Finally, another interesting model is Vector Quantised-VAE (VQ-VAE) (Van Den Oord et al., 2017); in this case, the encoder outputs discrete, rather than continuous, codes, and the prior is learned rather than static.
@@ 3.1.3. Applications
VAEs can be used for semi-supervised classification to provide an auxiliary objective, improving the data efficiency (Kingma et al., 2014; Maaløe et al., 2016); to perform iterative reasoning about objects in a scene (Eslami et al., 2016); to model the latent dynamics of an environment (Watter et al., 2015).
Of course, VAEs have also been used to generate synthetic data, including for conditional generation. For example, a layered foreground-background generative model can be used to generate images based on both the latent representation and a representation of the attributes (Yan et al., 2016). In (Gómez-Bombarelli et al., 2018) the latent space of a VAE is trained on chemical structures by means of gradient-based optimization toward certain properties (see Section 3.7). AAEs have also been applied to the same problem (Kadurin et al., 2017). Finally, another interesting application of VAE is Deep Recurrent Attentive Writer (DRAW) (Gregor et al., 2015). DRAW constructs scenes in an iterative way, by accumulating changes emitted by the decoder (then given to the encoder in input). This allows for iterative self-corrections and a more natural form of image construction. RNNs and attention mechanism are used to consider previous generations and to decide at each time step where to focus attention, respectively.
@@ 3.1.4. Critical Discussion
Models based on VAEs can be considered as an example of exploratory creativity. The latent space is learned with the goal of representing data in the most accurate way. The random sampling performed during generation is therefore an exploration of that space: regions not seen during training can be reached as well, even though they can lead to poor generation (Aneja et al., 2021) and some more complex variants may be needed, as discussed. On the other hand, there is no guarantee that the results will be valuable, novel, or surprising. There is no guarantee that the generation from random sampling is of good quality or diverse from training data. Indeed, given their characteristics, VAEs discourage novelty in a sense. In particular, diversity could be achieved in theory using VAEs and gradient-based optimization techniques, such as those presented in (Gómez-Bombarelli et al., 2018), with novelty and surprise as target properties. We will discuss these aspects in Section 3.8 in more detail.
@@ 3.2.1. Core Concepts
A Generative Adversarial Network (Goodfellow et al., 2014) is an architecture composed by two networks: a generative model and a discriminative model. The latter learns to distinguish between real samples and samples generated by the former. In parallel, the former learns to produce samples from random noise vector such that they are recognized as real by the latter. This competition drives both models to improve their methods, until the generated samples are indistinguishable from the original ones.
@@ 3.2.2. Examples of Models
In StyleGAN (Karras et al., 2019; Karras et al., 2020), the generator architecture is re-designed in order to control the image synthesis process. The style of the image is adjusted at each layer based on the latent code (the specific intermediate code to control each layer is provided by a non-linear mapping network). This allows for automatic separation of high-level attributes from stochastic variations in the generated images. It also allows for mixing regularization, where two latent codes are used alternatively to guide the generator. StyleGAN-V (Skorokhodov et al., 2022) builds on top of it to learn to produce videos by only using a few frames of it. To generate longer and more realistic motions, a two-stage approach can be used as well: first, a low-resolution generator is adversarially trained on long sequences; then, a high-resolution generator transforms a portion of the produced, low-resolution video in a high-resolution one (Brooks et al., 2022).
@@
Finally, it is also worth mentioning variants that adapt GANs to sequential tasks (e.g., text generation). Since GANs require the generator to be differentiable, they cannot generate discrete data (Goodfellow, 2017). However, several techniques have been proposed to avoid this problem. One possibility is to transform the discrete generation into a continuous one. Music can be processed like an image by considering its waveform (as in WaveGAN (Donahue et al., 2019) and GANSynth (Engel et al., 2019)) or its musical score composed of tracks and bars (as in MuseGAN (Dong et al., 2018)). Music in a desired style can be obtained through conditional inputs. Another possibility is to consider a soft-argmax function as an approximation of the inference for each step (Zhang et al., 2016). TextGAN (Zhang et al., 2017) uses it together with feature matching to learn the production of sentences. In place of the discriminative signal, it uses the difference between the latent feature distributions of real and of synthetic sentences learned by the discriminator. Another solution is to transform the GAN into a Reinforcement Learning (RL) framework, as in Sequential GAN (SeqGAN) (Yu et al., 2017). The generative model is the agent; the tokens generated so far from the state; the selection of the next token to be generated is the action to be performed; and the discriminative signal is the reward. The REINFORCE algorithm (Williams, 1992) can then be used to adversarially train the generative model. Other policy gradient methods can be used as well (Fedus et al., 2018). The advantage of using RL to adapt GANs to sequential tasks is that the reward function can be composed by more signals: additional objectives can be used as well to embed particular behaviors in the generator. As suggested by Objective-Reinforced GAN (ORGAN) (Guimaraes et al., 2017), qualitative heuristics - like tonality and ratio of steps - can be used to improve music generation. On the other hand, the learning signal (i.e., the reward) might be very sparse. A way to solve this issue is to use inverse RL (Ziebart et al., 2008). For example, the authors of (Shi et al., 2018) use inverse RL to learn a reward function able to associate positive rewards to real state-action pairs, and non-positive rewards to synthetic state-action pairs. Notably, this can help solve mode collapse too. Another variant is LeakGAN (Guo et al., 2018). Here, a hierarchical generator composed by a Manager and a Worker is used. The Worker produces a sentence conditioned by a goal vector provided by the Manager. The Worker and the discriminative model are trained following SeqGAN; the Manager is trained to predict goal vectors that lead to the identification of advantageous directions in the discriminative feature space. More specifically, the Manager receives a feature vector from the discriminator, i.e., its last convolutional layer, at each generated token. By means of this leaked information and the hierarchical architecture, LeakGAN produces longer and higher-quality texts. Finally, another possibility is to use Gumbel-softmax relaxation (Jang et al., 2017; Maddison et al., 2017), as in Relational GAN (RelGAN) (Nie et al., 2019). Controlled TExt generation Relational Memory GAN (CTERM-GAN) (Betti et al., 2020) builds on the latter by also conditioning the generator on an external embedding input. In addition, it uses both a syntactic discriminator to predict whether a sentence is correct and a semantic discriminator to infer if a sentence is coherent with the external input.

### v6

@@ 1. Introduction
The connection between creativity and machines is as old as computer science itself. In 1842, Lady Lovelace, an English mathematician and writer, who is recognized by many as the first computer programmer, issued what is now known as “Lovelace’s Objection” (Turing, 1950). She stated that the Analytical Engine (the digital programmable machine proposed by Charles Babbage (Babbage, 1864)) “has no pretensions to originate anything. It can do whatever we know how to order it to perform” (Menabrea and Lovelace, 1843). Indeed, in the centuries that followed, numerous projects and studies have been undertaken with the aim of designing machines capable of “originating something” (Cohen, 1988; Cope, 1989; Hofstadter and Mitchell, 1994; Langley et al., 1987; Meehan, 1977; Racter, 1984). We have witnessed the emergence of a specialized field in computer science, namely Computational Creativity (Cardoso et al., 2009), which concerns the study of the relationship between creativity and artificial systems (Colton and Wiggins, 2012; Wiggins, 2006).
@@
Related surveys. We now provide an overview of other surveys in areas related to the present work. For readers interested in a survey on deep generative models, we recommend (Bond-Taylor et al., 2021; Harshvardhan et al., 2020); for an analysis of the state of the art in evaluation in computational creativity, (Lamb et al., 2018) is an essential reading; for a review on AI and creativity in general, we recommend (Rowe and Partridge, 1993); for a practical view of generative deep learning, we suggest (Foster, 2019); finally, for an in-depth examination of artistic AI works (also in human-computer co-creations (Guzdial and Riedl, 2019)), (Miller, 2019) is a very comprehensive source of information.
@@ 2. Defining Creativity
For this reason, we consider Boden’s three criteria for studying machine creativity, defined as “the ability to come up with ideas or artifacts that are new, surprising and valuable (Boden, 2003). In particular, value refers to utility, performance, and attractiveness (Maher, 2010); it is related with both the quality of the production, and its acceptance by the society. Novelty refers to the dissimilarity between the produced artifact and other examples in its class (Ritchie, 2007). Finally, surprise refers to the degree to which a stimulus disagrees with expectation (Berlyne, 1971). We underpin our analysis on Boden’s three criteria since they have been widely adopted. Boden also suggests that three forms of creativity can be identified (Boden, 2003) to describe how a novel and surprising product is obtained. The three forms of creativity are combinatorial, exploratory and transformational, ordered by increasing rarity and produced surprise. Combinatorial creativity is about making unfamiliar combinations of familiar ideas, e.g., analogies in textual forms or collages in the visual arts. Exploratory creativity involves the exploration of the conceptual space defined by the cultural context considered, e.g., inventing a new type of cut for fries. Transformational creativity involves changing that space in a way that allows new and previously inconceivable thoughts to become possible, as it has been for free verse in poetry or abstract painting in art.
Finally, it is worth noting that Boden also identified four different questions that emerge when studying computational creativity. These are referred to as Lovelace questions, because many people would respond to them by using Lovelace’s objection. The first question is whether computational ideas can help us understand how human creativity works. The second is whether computers could ever do things that at least appear to be creative. The third is whether a computer could ever appear to recognize creativity. Finally, the fourth is whether computers themselves could ever really be creative, i.e., with the originality of the products only deriving from the machine itself (Boden, 2003). While the first one is studied in Boden’s work, we will provide the reader with an overview of the techniques that can possibly be used to answer the second (Section 3) and the third (Section 4). With respect to the fourth, Boden states that “it is not a scientific question as the others are, but in part a philosophical worry about “meaning” and in part a disguised request for a moral political decision”. We agree with this position, but we hope that our survey will provide the reader with elements for answering the fourth one as well.
@@ 3. Generative Models
At first glance, this definition appears to be incompatible with those presented in Section 2. Indeed, mimicry is the opposite of novelty. However, what a generative model should aim at mimicking is the underlying distribution representative of the artifacts, and not the specific artifacts themselves; in other words, it should aim at learning the conceptual space defined by the cultural context considered. A generative model can be said to exhibit combinatorial creativity if it can sample new and valuable works that are combinations of real data and exploratory creativity if the works actually differ from real ones. Vice versa, transformational creativity emerges if and only if the distribution for sampling diverges in some way from the underlying one (e.g., due to a different training process, by altering the distribution after learning, or by changing the sampling technique). In summary, what matters is how the space of solutions is learned and how artifacts are sampled from it.
In this section, we aim at studying the level of creativity of existing generative deep learning models. Following the discussion above, we analyze how the models learn their spaces of solutions and how the observations are generated from them. A new generative deep learning taxonomy is then introduced based on the different training and sampling techniques at the basis of each method. Figure 1 provides a summary of the seven generative classes considered in this survey. Since our focus is on machine creativity, we do not discuss the implementation details of each class of methods. We instead present the core concepts at the basis of each class; some relevant examples of models; potential applications; and a critical discussion evaluating the level of machine creativity considering the definitions above.
@@ 3.1.1. Core Concepts
A Variational Auto-Encoder (VAE) (Kingma and Welling, 2014; Rezende et al., 2014) is a learning architecture composed of two models: an encoder (or recognition model) and a decoder (or generative model). The former compresses high-dimensional input data into a latent space, i.e., a lower-dimensional space whose features are not directly observable, yet provide a meaningful representation. The latter decompresses the representation vector back to the original domain (Foster, 2019). Classic autoencoders directly learn to represent each input in a latent representation vector. Conversely, VAEs learn a (Gaussian) distribution over the possible values of the latent representation, i.e., the encoder learns the mean and the (log of the) variance of the distribution.
@@
The mathematical derivation of the whole loss has its roots in variational inference (Jordan et al., 1999). Indeed, VAEs can be seen as an efficient and stochastic variational inference method, in which neural networks (NNs) and stochastic gradient descent are used to learn an approximation (i.e., the encoder) of the true posterior (Ganguly and Earp, 2021). In VAEs, similar high-dimensional data are mapped to close distributions. This makes it possible to sample a random point $\mathbf{z}$ from the latent space, and still obtain a comprehensible reconstruction (Foster, 2019). On the other hand, VAE tends to produce blurred images (Zhao et al., 2017). It may also happen that high-density regions under the prior have a low density under the approximate posterior, i.e., these regions are not decoded to data-like samples (Aneja et al., 2021). Finally, the objective can lead to overly simplified representations without using the entire capacity, obtaining only a sub-optimal generative model (Burda et al., 2016).
@@ 3.1.2. Examples of Models
Several models based on VAEs have been proposed (Kingma and Welling, 2019) in recent years. In the following, we focus on those relevant to our discussion on machine creativity. In $\beta$-VAE (Higgins et al., 2017), a parameter $\beta$ is used to scale the magnitude of the regularization loss, which allows a better disentanglement of the latent space (Burgess et al., 2018). Another example is VAE-GAN (Larsen et al., 2016), which merges VAE and Generative Adversarial Networks (GAN; see Section 3.2) (Goodfellow et al., 2014). This is done by treating the decoder as the generator of the GAN, thus training it by means of the GAN loss function. This leads to the generation of substantially less blurred images. Similarly, Adversarially Learned Inference (ALI) (Dumoulin et al., 2017) merges VAE and GAN by asking the discriminator to distinguish between pairs of real data (and their latent representations) and pairs of sampled representations and synthetic data. Instead, Adversarial Autoencoders (AAE) (Makhzani et al., 2016) substitute the regularization loss with a discriminative signal, where the discriminator has to distinguish between random latent samples and encoded latent vectors. Another way to address the problem of “sample blurriness” is with PixelVAE (Gulrajani et al., 2017), where the autoregressive PixelCNN (Van Den Oord et al., 2016b, c) is used as the decoder. In (Bowman et al., 2016), to deal with sequential data such as texts, where generation requires more steps, the encoder learns to produce a latent representation of a sentence, while the recurrent neural network RNN-based decoder learns to reproduce it word after word. However, VAE can also generate text by means of convolution and deconvolution (Semeniuta et al., 2017). To solve the problem of low-density regions, the authors of (Aneja et al., 2021) propose an energy-based model called noise contrastive prior (NCP), trained by contrasting samples from the aggregate posterior to samples from a base prior. Finally, another interesting model is Vector Quantised-VAE (VQ-VAE) (Van Den Oord et al., 2017); in this case, the encoder outputs discrete, rather than continuous, codes, and the prior is learned rather than static.
@@ 3.1.3. Applications
VAEs can be used for semi-supervised classification to provide an auxiliary objective, improving the data efficiency (Kingma et al., 2014; Maaløe et al., 2016); to perform iterative reasoning about objects in a scene (Eslami et al., 2016); to model the latent dynamics of an environment (Watter et al., 2015). Of course, VAEs have also been used to generate synthetic data, including for conditional generation. For example, a layered foreground-background generative model can be used to generate images based on both the latent representation and a representation of the attributes (Yan et al., 2016). In (Gómez-Bombarelli et al., 2018) the latent space of a VAE is trained on chemical structures by means of gradient-based optimization toward certain properties (see Section 3.7). AAEs have also been applied to the same problem (Kadurin et al., 2017). Finally, another interesting application of VAE is Deep Recurrent Attentive Writer (DRAW) (Gregor et al., 2015). DRAW constructs scenes in an iterative way, by accumulating changes emitted by the decoder (then given to the encoder in input). This allows for iterative self-corrections and a more natural form of image construction. RNNs and attention mechanism are used to consider previous generations and to decide at each time step where to focus attention, respectively.
@@ 3.1.4. Critical Discussion
Models based on VAEs can be considered as an example of exploratory creativity. The latent space is learned with the goal of representing data in the most accurate way. The random sampling performed during generation is therefore an exploration of that space: regions not seen during training can be reached as well, even though they can lead to poor generation (Aneja et al., 2021) and some more complex variants may be needed, as discussed. On the other hand, there is no guarantee that the results will be valuable, novel, or surprising. There is no guarantee that the generation from random sampling is of good quality or diverse from training data. Indeed, given their characteristics, VAEs discourage novelty in a sense. In particular, diversity could be achieved in theory using VAEs and gradient-based optimization techniques, such as those presented in (Gómez-Bombarelli et al., 2018), with novelty and surprise as target properties. We will discuss these aspects in Section 3.8.
@@ 3.2.1. Core Concepts
A Generative Adversarial Network (Goodfellow et al., 2014) is an architecture composed by two networks: a generative model and a discriminative model. The latter learns to distinguish between real samples and samples generated by the former. In parallel, the former learns to produce samples from random noise vectors such that they are recognized as real by the latter. This competition drives both models to improve their methods until the generated samples are indistinguishable from the original ones.
@@ 3.2.2. Examples of Models
In StyleGAN (Karras et al., 2019; Karras et al., 2020), the generator architecture is re-designed in order to control the image synthesis process. The style of the image is adjusted at each layer based on the latent code (the specific intermediate code to control each layer is provided by a non-linear mapping network). This allows for the automatic separation of high-level attributes from stochastic variations in the generated images. It also allows for mixing regularization, where two latent codes are used alternatively to guide the generator. StyleGAN-V (Skorokhodov et al., 2022) builds on top of it to learn to produce videos by only using a few frames of it. To generate longer and more realistic motions, a two-stage approach can be used as well: first, a low-resolution generator is adversarially trained on long sequences; then, a high-resolution generator transforms a portion of the produced, low-resolution video in a high-resolution one (Brooks et al., 2022).
@@
Finally, it is also worth mentioning variants that adapt GANs to sequential tasks (e.g., text generation). Since GANs require the generator to be differentiable, they cannot generate discrete data (Goodfellow, 2017). However, several techniques have been proposed to avoid this problem. One possibility is to transform the discrete generation into a continuous one. Music can be processed like an image by considering its waveform (as in WaveGAN (Donahue et al., 2019) and GANSynth (Engel et al., 2019)) or its musical score composed of tracks and bars (as in MuseGAN (Dong et al., 2018)). Music in a desired style can be obtained through conditional inputs. Another possibility is to consider a soft-argmax function as an approximation of the inference for each step (Zhang et al., 2016). TextGAN (Zhang et al., 2017) uses it together with feature matching to learn the production of sentences. In place of the discriminative signal, it uses the difference between the latent feature distributions of real and synthetic sentences learned by the discriminator. Another solution is to transform the GAN into a Reinforcement Learning (RL) framework, as in Sequential GAN (SeqGAN) (Yu et al., 2017). The generative model is the agent; the tokens generated so far form the state; the selection of the next token to be generated is the action to be performed; and the discriminative signal is the reward. The REINFORCE algorithm (Williams, 1992) can then be used to adversarially train the generative model. Other policy gradient methods can be used as well (Fedus et al., 2018). On the other hand, the learning signal (i.e., the reward) might be very sparse. A way to solve this issue is to use inverse RL (Ziebart et al., 2008). For example, the authors of (Shi et al., 2018) use inverse RL to learn a reward function able to associate positive rewards to real state-action pairs, and non-positive rewards to synthetic state-action pairs. Notably, this can help solve mode collapse too. Another variant is LeakGAN (Guo et al., 2018). Here, a hierarchical generator composed of a Manager and a Worker is used. The Worker produces a sentence conditioned by a goal vector provided by the Manager. The Worker and the discriminative model are trained following SeqGAN; the Manager is trained to predict goal vectors that lead to the identification of advantageous directions in the discriminative feature space. More specifically, the Manager receives a feature vector from the discriminator, i.e., its last convolutional layer, at each generated token. By means of this leaked information and the hierarchical architecture, LeakGAN produces longer and higher-quality texts. Finally, another possibility is to use Gumbel-softmax relaxation (Jang et al., 2017; Maddison et al., 2017), as in Relational GAN (RelGAN) (Nie et al., 2019). Controlled TExt generation Relational Memory GAN (CTERM-GAN) (Betti et al., 2020) builds on the latter by also conditioning the generator on an external embedding input. In addition, it uses both a syntactic discriminator to predict whether a sentence is correct and a semantic discriminator to infer if a sentence is coherent with the external input.